OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!!

OpenStudy (abbycross167):

Please help me to simplify this rational expression -- (8y) (7x) ------ * ------- (14x^2) (8z^3)

OpenStudy (phi):

the idea is to concentrate first on just the numbers if you had \[ \frac{8 \cdot 7 }{14 \cdot 8} \] could you simplify this ? If I were doing it, I would write it as \[ \frac{8}{8} \cdot \frac{7}{14} \]

OpenStudy (abbycross167):

May I ask why you would write it like that?

OpenStudy (phi):

as you know, we can reorder multiplies and divisions and 8/8 is easy and 7/14 is almost as easy it's a lot easier than doing 7*8= 56 and 14*8= 112 and simplifying 56/112

OpenStudy (abbycross167):

Oh okay I understand now!

OpenStudy (abbycross167):

So 7/14 = 1/2 What about the x^2?

OpenStudy (abbycross167):

Ok so would I be left with just one x?

OpenStudy (phi):

now z's. but you only have z^3 in the bottom. so we are done: \[ \frac{xy}{2z^3}\]

OpenStudy (phi):

oh wait, the x's should have simplified to 1/x so the answer is \[ \frac{y}{2xz^3} \]

OpenStudy (phi):

the \[ \frac{x}{x^2} = \frac{x}{x} \cdot \frac{1}{x} =\frac{1}{x}\]

OpenStudy (abbycross167):

Sir I think some of your replies are being deleted... I can't see them anymore

OpenStudy (phi):

yes, I deleted the posts that had mistakes. Less confusing (I hope)

OpenStudy (abbycross167):

Oh okay, I just wanted to make sure I wouldn't miss a step... Thank you very much for your help!! I appreciate it!!